At what velocity did you travel 250.0 m after 4.0 min if you started from rest?


1
Expert's answer
2021-10-26T10:01:11-0400

We can find the final velocity from the kinematic equation:


"d=\dfrac{1}{2}(v_0+v)t,""v=\dfrac{2d}{t}=\dfrac{2\times250\ m}{4\ min\times\dfrac{60\ s}{1\ min}}=2.1\ \dfrac{m}{s}."